A thermochromic plastic chip included in a shipping container for beef undergoes an irreversible color change if the storage temperature exceeds 28°F.
What is this temperature on the Celsius scale?
express your answer in two sig figs
What is this temperature on the Kelvin scale?
express your answer in four sig figs
Answer #1

(C/100)=(F-32)/180

C is the temperature in degree celsius

F is the temeprature in Fahrenheit

C/100 = (28-32)/180

C=-2.20C

Temperature in kelvin=temperature in 0C+273.15

Temperature in kelvin=-2.2+273.15 =270.95 K=271.0 K